RFBG Commissioning with Adaptive Secondary - 27 April 2010 UT

Night 19:

Observer: JHeidt (LBTO), PHibon (LBTO), DMiller (Tucson)
Telescope Operator: SAllansen (LBTO), DHuerta (LBTO)
Telescope Support: JUrban (LBTO)
Instrument Support: JMorris (LBTO)
Adaptive Secondary Operator: MXompero (LBTO)

Summary

Night hampered by cirrus, but offered subarcsec seeing

More than 3hrs at beginning of night lost due to a wide variety of (humanmade) problems (cf report by TO)

Moved back to some cfg-files of A03 of March 29

Got a good set of transformation data

Focus measurements for IRTC

Marco did some tests with an M5 star at the very end of the night

Preparations

IRTC was successfully moved from LDG to RFBG during day.

Single frame names refer to: irtc.20100427.xxxxxx.fits

Cube names refer to: irtc.20100427_00xxx_cube.fit

Plan for the night: Collimation issues, transformation data, field aberration data, Z6 as function of elevation (for Marco) if seeing is bad

Steve is his first night at the telescope, David is going to train him

Details

Open

02:02 Hexapod in simulation mode, called Paul, back to normal life, homing secondary, OK

02:11 open, lots of cirrus around and some thicker clouds to the SE, T_out = 5.5C, T_mirror = 5.6C, Hum = 35%, low wind (2-3m/s)

02:12 Marco has an issue with the secondary (he didn't specify what) but should by now not affect us

02:15 Some delays, telescope not yet ready, see TO's report for details

Initial Pointing + Collimation

02:50 pointto Pollux TelM=Track, gs=0, PAM-, doesn't work, could not authorize IRTC to RFBG

02:25 Called Dave: need to epar LBTtools and set camera to LUC2, epar IRTC and set telside to right

03:20 pointto Pollux TelM=Track, gs=0, PAM-, preset failed but this time for a different reason, UMAC communication has been disabled, tertiary mirror needs to be set in simulation mode, called Dan, had to be initialized

03:44 pointto Pollux TelM=Track, gs=0, PAM-, preset works, but no star

03:45 pointto Regulus TelM=Track, gs=0, PAM-, works, EL=67, too low

03:49 pointto M5_0425, TelM=track, gs=0, PAM-, EL=67, no star, ptspiral, no star?, more clouds coming in

04:03 pointto Regulus TelM=Track, gs=0, PAM-, EL=65, ptspiral, no star...

04:35 David went up, tertiary swing arm not in place, apparently John didn't finish his job in the afternoon completely...can not blame him, it the responsibility of the lead observer (Jochen) to make sure that everything is OK. JMH: You can blame me a little bit.

04:40 David deploys the tertiary swing arm, dynamic balancing

05:00 pointto M5_1332, TelM=track, gs=0, PAM-, EL=78, no star, ptspiral, no success

05:00 pointto M5_1332, TelM=acquire, gs=0, PAM-, EL=78, no star, ptspiral again, no success

05:35 David went up again, tertiary mirror is still not in the right position, need to call Paul, tertiary was re-initialized

06:06 pointto M5_1332, TelM=track, gs=0, PAM-, EL=83, failed several times (got error message)

06:16 pointto M5_1332, TelM=track, gs=0, PAM-, EL=83, preset OK, but no star

06:17 pointto M5_1332, TelM=acquire, gs=0, PAM-, EL=83, preset OK, no star, ptspiral, YES, star found

06:22 pointto M5_1332 (high-el), TelM=track, gs=0, PAM-, EL=83, CA=-77, IE -5, remove coma, setting M2 RX=-540, RY=-20 and CA=-64, IE=30

06:35 pointto M5_0519 (low-el), TelM=track, gs=0, PAM-, EL=30, IA=-387, IE=35

06:42 pointto BS9148, TelM=active, gs=0, EL=71, converged below 200 rms, guider reports ~0.5" seeing

Transformation data

06:54 pointto StoneLagw, TelM=active, gs=34, PA=180, collimate, seeing on guider 0.6-0.7"

07:03 offset 0 30, takeseq 2 30 070428 cube 1, backgr. image, offset 0 -30

07:10 cl < tran_StoneL_mix1.cl

Clouds/cirrus passing from time to time

PA = 0

transtar (35) 071129 cube 2
transtar (31) 071336 cube 3
transtar (21) 071543 cube 4
transtar (19) 071746 cube 5
transtar (15) 071952 cube 6

PA = 180

transtar (23) 072232 cube 7
transtar (17) 072438 cube 8
transtar (20) 072644 cube 9
transtar (18) 072850 cube 10

PA = 45

transtar (35) 073139 cube 11
transtar (31) 073344 cube 12
transtar (13) 073552 cube 13
transtar (24) 073758 cube 14
transtar (30) 074019 cube 15, clouds, no gs found, only be used as backgr.

PA = 210

transtar (12) 074252 cube 16, potentially hampered by clouds
transtar (22) 074458 cube 17
transtar (26) 074702 cube 18
transtar (34) 074906 cube 19

07:52 Stopped, too many clouds

Focus IRTC

07:53 pointto BS9162 TelM=active gs=6 PA=indef, guider reports 0.5" and below!

08:00 offset 0 30 takeseq 1.5 30 080206 cube 20, offset 0 -30

08:04 cl < focus_seq.cl, aborted forgot to turn off WFS

08:09 pointto BS9162 TelM=active gs=6 PA=indef

08:16 cl < focus_seq.cl, x=0.04, y=58.55, z=19.27, delta=1.73

Images 081659 cube 23 ... 082702 cube 30

08:09 pointto BS9162 TelM=active gs=6 PA=indef

08:31 cl < focus_seq.cl

Images 083204 cube 31 ... 084204 cube 38

08:44 cl < focus_seq.cl

Images 084517 cube 39 ... 085519 cube 46

Transformation data

09:05pointto StoneMagw, TelM=active, gs=55, PA=indef, collimate took quite a while, seeing on guider

09:13 offset 0 30 takeseq 2 30 091357 cube 47, for backgr., offset 0 -30

09:17 cl < home$tran_StoneM_mix1.cl

PA = 0

transtar (39) 091903 cube 48
transtar (72) 092115 cube 49
transtar (52) 092323 cube 50

PA = 90

transtar (55) 092546 cube 51
transtar (64) 092756 cube 52
transtar (58) 093003 cube 53
transtar (85) 093216 cube 54
transtar (79) 093424 cube 55

PA = 180

transtar (55) 093645 cube 56

09:37 Stopped, collimation diverges!

09:38 pointto StoneMagw TelM=active gs=0, collimate, OK, guider reports ~0.5-0.6"

09:43 pointto StoneMagw TelM=active gs=70, PA=indef, EL=55, does not collimate properly

09:56 pointto StoneMagw TelM=active gs=0, collimate

09:59 On request by Doug, David replaces config A04-file by the one for A03 of March 29 (AIP_IRTC_R.cfg)

10:02 pointto StoneMagw TelM=active gs=70, PA=indef, EL=55, no improvement (even slightly more worse)

10:12 On request by Doug, David replaces another config A04-file by the one for A03 of March 29 (AIP_IRTC_RM_SUB11.cfg)

10:02 pointto StoneMagw TelM=active gs=70, PA=indef, collimates well,

10:17 pointto StoneMagw TelM=active gs=70, PA=80, collimates also well

10:25 cl < home$tran_StoneM_mix1.cl, seeing 0.5 - 0.7" on guider during script

PA = 0

transtar (39) 102554 cube 57
transtar (72) 102809 cube 58
transtar (52) 103015 cube 59

PA = 90

transtar (55) 103238 cube 60
transtar (64) 103448 cube 61
transtar (58) 103656 cube 62
transtar (85) 103908 cube 63
transtar (79) 104044 cube 64, rotator stuck, not useful, but as backgr

PA = 180

transtar (55) 104220 cube 65, rotator stuck, not useful, but as backgr., rotator reset by David
transtar (76) 104356 cube 66, rotator stuck, not useful
transtar (78) 104616 cube 67, OK
transtar (60) 104826 cube 68
transtar (54) 105034 cube 69

PA = 0

transtar (43) 105313 cube 70
transtar (35) 105518 cube 71
transtar (13) 105727 cube 72

PA = 90

transtar (74) 105948 cube 73
transtar (63) 110157 cube 74
transtar (49) 110405 cube 75
transtar (55) 110613 cube 76

PA = 180

transtar (50) 110834 cube 77
transtar (46) 111045 cube 78, guide star too faint, not useful
transtar (32) 111255 cube 79
transtar (31) 111432 cube 80, rotator got stuck, not useful except for backgr.
transtar (16) 111638 cube 81, guide star too faint, not useful

10:18 cl < home$tran_StoneM_mix2.cl, seeing 0.5 - 0.7" on guider during script

PA = 45

transtar (74) 111915 cube 82
transtar (67) 112122 cube 83
transtar (72) 112337 cube 84
transtar (58) 112544 cube 85
transtar (44) 112750 cube 86

PA = 135

transtar (41) 113015 cube 87
transtar (54) 113226 cube 88
transtar (60) 113434 cube 89
transtar (70) 113640 cube 90
transtar (76) 113848 cube 91

PA = 225

transtar (20) 114024 cube 92, rotator stuck, not useful except backgr.
transtar (10) 114200 cube 93, rotator stuck, not useful, except backgr., David reset rotator
transtar (27) 114420 cube 94
transtar (50) 114626 cube 95
transtar (60) 114832 cube 96

PA = 315

transtar (58) 115129 cube 97
transtar (44) 115336 cube 98
transtar (31) 115548 cube 99
transtar (20) 115759 cube 100
transtar (11) 120014 cube 101

Its getting bright, the last one or two measurements are probably not good enough, check!

Marco's doing some tests via Italy

12:02 pointto M5_0619 TelM=active, gs=1, PA=indef, OK

Determine interaction matrix between AGw and mirror on sky in order to check orientation of the secondary wrt to the WFS

Close

12:30 Closed, still cirrus, T_out = 3.7, T_Mirror = 3.9, Hum = 40%

-- JochenHeidt - 27 Apr 2010
Topic revision: r7 - 29 Apr 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback